zolo:

What is the value of x? Four-fifths x minus StartFraction 1 Over 10 Endfraction = StartFraction 3 Over 10 EndFraction

jhonyy9:

\[\frac{ 4 }{ 5 }x - \frac{ 1 }{ 10 } = \frac{ 3 }{ 10 }\] so like a first step what is the common denominator for 5 and 10 on the left side ?